Three numbers are so related that the second number is 2 more than the first number and third number is 4 more than the first. Find the numbers if the sum of their squares is 56 more than three times the square of the smallest number.
Answer by sachi(548)

let the 1st no =x
2nd no =x+2
& 3rd no =x+4
the sum of their squares is 56 more than three times the square of the smallest number.
so x^2+(x+2)^2+(x+4)^2=56+3*x^2
or x^2+x^2+4x+4+x^2+8x+16=56+3x^2
or3x^2+12x+20=56+3x^2
or 12x=56-20=36
or x=36/12=3
so the nos are 3,5,7 ans